OpenStudy (anonymous):

Use an angle sum identity to verify the identity cos 2 0 = cos^2 0 -1 .

OpenStudy (reemii):

can you put parentheses? it's difficult to understand whaty ou mean..

OpenStudy (jdoe0001):

$$ cos(2\theta) = cos^2(\theta)-1? $$

OpenStudy (reemii):

what is "cos 2 0" ?

OpenStudy (anonymous):

im not honestly sure i am horriable with algebra its deff my weakest point

OpenStudy (anonymous):

cos 2θ = cos^2(θ) − 1/2 ---------------------------------- (cos x)^2 - (sin x)^2 = (cos x)^2 - 1/2 (sin x)^2 = 1/2 or sin x = +/- 1 / sqrt(2) = +/- sqrt(2) / 2 x = +/- pi/4. 3.(pi)/4, 5.(pi)/4, 7.(pi)/7 --------------------------------------… http://answers.yahoo.com/question/index?qid=20120326173305AA0BuSA

OpenStudy (reemii):

the real formula is \(\cos(2\theta) = 2\cos^2\theta - 1\).